# Peptide Classification from Statistical Analysis of Nanopore Translocation Experiments

### v1

@@ Abstract
Protein characterization using nanopore-based devices promises to be a breakthrough method in basic research, diagnostics, and analytics. Current research includes the use of machine learning to achieve this task. In this work, a comprehensive statistical analysis of nanopore current signals is performed and demonstrated to be sufficient for classifying up to 42 peptides with 70 % accuracy. Two sets of features, the statistical moments and the catch22 set, are compared both in their representations and after training small classifier neural networks. We demonstrate that complex features of the events, captured in both the catch22 set and the central moments, are key in classifying peptides with otherwise similar mean currents. These results highlight the efficacy of purely statistical analysis of nanopore data and suggest a path forward for more sophisticated classification techniques.
@@
Peptide characterization, Nanopore blockades, Machine learning
@@ 2.2 Model training
Table 1: Mean classification results after training with six different seeds for the model initialization.
@@
Moment decomposition $0.704(3)$ $0.708(3)$ $0.718(2)$ $0.705(3)$
Catch22 $0.715(3)$ $0.723(2)$ $0.728(3)$ $0.720(3)$
Table 1 show that the catch22 trained model achieved marginally better classification performance than the moment decomposition, although it still failed to discern all peptide ladder classes fully. The precision score, which indicates how often a peptide ladder was correctly predicted compared to the total number of times it was predicted, is slightly improved in the catch22 model, suggesting a smaller number of false predictions on average. The recall scores show a similar trend, with the catch22 trained models less likely to miss a true classification. Therefore, it is unsurprising that the $F_{1}$ Score, comprised of both precision and recall, also follows this trend. This relatively simple analysis method can classify the 42 peptide classes surprisingly well. However, in many of these cases, mean current alone can be used, as outlined in Behrends and Ensslen 25. Therefore, to highlight the specific role of the higher-order moments, the model capabilities of distinguishing between the six decapeptides is shown in Figure 4. These decapeptides are chosen as they have identical mass, resulting in a high degree of overlap in the event mean histogram, making prediction based on the mean alone infeasible. Specifically, the confusion matrix in Figure 4b shows that identification of the decapeptides can be achieved beyond the straightforward comparison of the mean current of the normalized events, highlighting that additional information lies in higher-order moments of the signal, and supporting the use of more sophisticated classification methods. Note that the histograms in Figure 4a do not represent the optimally achievable identification of the mean current. Rather, the event detection explicitly includes parts of the rising and falling phase in each event, which broadens and possibly skews all peaks. This can be accounted for by employing additional steps in the event preprocessing 21. For this task however, we explicitly include the rise and fall of each event based on the assumption that they might also contribute to the features in both feature sets.
@@ 2.3 SHAP Model Explanation
In a standard SHAP figure, the SHAP value indicates whether the variable increased (positive) or decreased (negative) the model output. The color of each point indicates whether a large (red) or small (blue) scalar value of that feature, the feature value, led to the change. For example, a small feature value with a large SHAP score would indicate that small scalar values for that feature were correlated with larger model outputs. To study the impact of the features with respect to their scalar values, the absolute SHAP values are shown. Examining Figure 5, several interesting outcomes are evident. First, some features shared between the data sets have strongly different SHAP scores in the two models. This is most prevalent in the relative mean current value: whereas a higher value of the relative mean current plays a more important role in the final prediction of the model trained on the moments decomposition dataset, the model trained on the catch22 dataset behaves the opposite, with smaller values having more impact to the model prediction. This behavior could be the result of correlation between features, something that is not taken into account in this SHAP analysis. Specifically, while the catch22 dataset contains features that were explicitly chosen for their lack of dependence to one another, the same cannot be said for the moment decomposition feature set. In both cases, the dominance of the mean is still prevalent in the predictions. The catch22-trained models also seemed to depend greatly on the autocorrelation time as estimated by computing the autocorrelation function of an event, indicating that dynamic properties also have an impact on the model prediction. This was completely missed in the statistical moment analysis but might be key to further improvements in predictive performance.
@@ 3 Conclusions
We have demonstrated that the moment decomposition of the time series data generated in nanopore experiments, either through the moments of the distributions formed by the events or more sophisticated analysis with the catch22 feature set, can be used to construct feature vectors for input into classification models. Based on the non-Gaussian form of the histograms, these features capture information present in the higher order moments of the current signal distributions. We showed that using small neural networks, classification accuracies of 70.4 % and 71.5 % can be achieved for the moments and catch22, respectively, with up to 42 peptide classes. These results were emphasized by showing the performance on the decapeptides where single event distinction is hindered by overlaps in the relative current mean histogram. We further found through SHAP analysis that the signal’s dynamic properties yielded additional classification performance in the catch22 dataset and should be considered in more complex feature decomposition approaches. This simple analysis is computationally efficient and capable of previously unseen classification performance on many classes. Our results highlight the role machine-learning classification can play in classification of large numbers of peptides that may one day be used in a clinical setting. Further, we demonstrate that preparative separation of peptides in experimental settings may not be necessary, as our models can differentiate those with highly overlapping mean blockade currents.

### v2

@@ Abstract
Protein characterization using nanopore-based devices promises to be a breakthrough method in basic research, diagnostics, and analytics. Current research includes the use of machine learning to achieve this task. In this work, a comprehensive statistical analysis of nanopore current signals is performed and demonstrated to be sufficient for classifying up to 42 peptides with over 70 % accuracy. Two sets of features, the statistical moments and the catch22 set, are compared both in their representations and after training small classifier neural networks. We demonstrate that complex features of the events, captured in both the catch22 set and the central moments, are key in classifying peptides with otherwise similar mean currents. These results highlight the efficacy of purely statistical analysis of nanopore data and suggest a path forward for more sophisticated classification techniques.
@@
Peptide characterization, Nanopore translocation, Machine learning
@@ 2.2 Model training
Table 1: Classification results of the dense neural networks after training on each feature set. The precision, recall and $F_{1}$-score computed by weighting each class metric by the number of available labels of that class.
@@
Moment decomposition $0.728$ $0.731$ $0.730$ $0.726$
Catch22 $0.736$ $0.737$ $0.736$ $0.733$
Table 1 show that the catch22 trained model achieved marginally better classification performance than the moment decomposition, although it still failed to discern all peptide ladder classes fully. The precision score, which indicates how often a peptide ladder was correctly predicted compared to the total number of times it was predicted, is slightly improved in the catch22 model, suggesting a smaller number of false predictions on average. The recall scores show a similar trend, with the catch22 trained models less likely to miss a true classification. Therefore, it is unsurprising that the $F_{1}$-score, comprised of both precision and recall, also follows this trend. This relatively simple analysis method can classify the 42 peptide classes surprisingly well. However, in many of these cases, mean current alone can be used, as outlined in Behrends and Ensslen 25. Therefore, to highlight the specific role of the higher-order moments, the model capabilities of distinguishing between the six decapeptides is shown in Figure 4. These decapeptides are chosen as they have identical mass, resulting in a high degree of overlap in the event mean histogram, making prediction based on the mean alone infeasible. Specifically, the confusion matrix in Figure 4b shows that identification of the decapeptides can be achieved beyond the straightforward comparison of the mean current of the normalized events, highlighting that additional information lies in higher-order moments of the signal, and supporting the use of more sophisticated classification methods. Note that the histograms in Figure 4a do not represent the optimally achievable identification of the mean current. Rather, the event detection explicitly includes parts of the rising and falling phase in each event, which broadens and possibly skews all peaks. This can be accounted for by employing additional steps in the event preprocessing 21. For this task however, we explicitly include the rise and fall of each event based on the assumption that they might also contribute to the features in both feature sets.
@@ 2.3 SHAP Model Explanation
In a standard SHAP figure, the SHAP value indicates whether the variable increased (positive) or decreased (negative) the model output. The color of each point indicates whether a large (red) or small (blue) scalar value of that feature, the feature value, led to the change. For example, a small feature value with a large SHAP score would indicate that small scalar values for that feature were correlated with larger model outputs. To study the impact of the features with respect to their scalar values, the absolute SHAP values are shown. Examining Figure 5, several interesting outcomes are evident. First, input features present in both datasets, such as the mean blocking current and the logarithmic dwell time, show similar importance characteristics. The catch22-trained models also seemed to depend greatly on the autocorrelation time as estimated by computing the autocorrelation function of an event, indicating that dynamic properties also have an impact on the model prediction. This was completely missed in the statistical moment analysis but might be key to further improvements in predictive performance.
Not taken into account are correlations between features, which was not captured by the SHAP analysis. However, the similarity in importance scores of the inputs present in both datasets suggests that the correlations of inputs are negligible.
@@ 3 Conclusions
We have demonstrated that the moment decomposition of the time series data generated in nanopore experiments, either through the moments of the distributions formed by the events or more sophisticated analysis with the catch22 feature set, can be used to construct feature vectors for input into classification models. Based on the non-Gaussian form of the histograms, these features capture information present in the higher order moments of the current signal distributions. We showed that using small neural networks, classification accuracies of 72.8 % and 73.6 % can be achieved for the moments and catch22, respectively, with up to 42 peptide classes. These results were emphasized by showing the performance on the decapeptides where single event distinction is hindered by overlaps in the relative current mean histogram. We further found through SHAP analysis that the signal’s dynamic properties yielded additional classification performance in the catch22 dataset and should be considered in more complex feature decomposition approaches. This simple analysis is computationally efficient and capable of previously unseen classification performance on many classes. Our results highlight the role machine-learning classification can play in classification of large numbers of peptides that may one day be used in a clinical setting. Further, we demonstrate that preparative separation of peptides in experimental settings may not be necessary, as our models can differentiate those with highly overlapping mean blockade currents.